# VulMCI : Code Splicing-based Pixel-row Oversampling for More Continuous Vulnerability Image Generation

### v1

@@ 3.1. Experiment Settings
3.2. Experimental Results
3.2.1. Experiments for Answering RQ1
VulMCI, the method proposed by us for vulnerability detection, demonstrates excellent performance across multiple evaluation metrics, achieving an F1 Score of 92.645% and an accuracy (ACC) of 89.157%. Figure 4 illustrates the changes in loss value and accuracy during the training of VulMCI over 200 epochs. It can be observed that the loss value gradually decreases and levels off around 0.11, while the training accuracy steadily increases and fluctuates around 92.8%. This provides evidence of the effectiveness of our model. To assess the impact of the concatenated line parameter , we conducted a parameter analysis experiment, and the results are presented in Table 3.
| k | FPR | FNR | Pr | Re | F1 | ACC |
| 1 | 21.429 | 9.23 | 92.221 | 90.77 | 91.49 | 87.559 |
| 2 | 20.476 | 8.677 | 92.583 | 91.323 | 91.949 | 88.217 |
| 3 | 20.595 | 8.464 | 92.559 | 91.536 | 92.044 | 88.342 |
| 4 | 17.738 | 9.315 | 93.468 | 90.685 | 92.055 | 88.468 |
| 5 | 18.69 | 8.847 | 93.174 | 91.153 | 92.152 | 88.562 |
| 6 | 20.476 | 8.592 | 92.589 | 91.408 | 91.995 | 88.28 |
| 7 | 19.048 | 8.805 | 93.056 | 91.195 | 92.116 | 88.499 |
We evaluated the effectiveness of concatenated lines by setting 7 values for in the analysis. As shown in Table 3, when ranges from 1 to 5, larger values of indicate more inserted concatenated lines, resulting in enhanced image continuity and gradually improving detection accuracy. Compared to (no inserted concatenated lines) with an accuracy (ACC) of 87.559%, the concatenation operation brings about a roughly 1% improvement. When , the detection performance fluctuates but overall remains at a high level.
3.2.2. Experiments for Answering RQ2
@@
In this section, we compare VulMCI with several vulnerability detection tools, including a commercial static vulnerability detection tool (i.e., Checkmarx(Checkmarx, 2021)), two open-source static analysis tools (i.e., FlawFinde(FlawFinder, 2021) and RATS(Rough Audit Tool for Security, 2021)), and four deep learning-based vulnerability detection methods (i.e., VulDeePecker(Li et al., 2018), VulCNN(Wu et al., 2022), Devign(Zhou et al., 2019), and SySeVR(Li et al., 2022)). Figure 5 presents a bar chart that intuitively illustrates the metrics of True Positive Rate (TPR), True Negative Rate (TNR), and Accuracy (ACC) for each of these methods.
@@
For the three previous static analysis tools, RATS, Checkmarx, and FlawFinder, their detection performance is not satisfactory, with accuracy all below 60%. One possible explanation is that they rely on human experts to define vulnerability rules, and as vulnerability types become more complex, human experts may not be able to comprehensively define all patterns, leading to higher false positives and false negatives.
@@
For the VulDeePecker and SySeVR methods, both employ program slicing on the dataset. They vectorize the slices and use them to train Bidirectional Recurrent Neural Networks (BRNN) for vulnerability detection. The performance disparity between these two systems stems from VulDeePecker solely considering vulnerabilities caused by API function calls and only accounting for data dependency information. On the other hand, SySeVR incorporates semantic information caused by both data and control dependencies. It supports a broader range of vulnerability syntax labels for matching vulnerabilities. Consequently, SySeVR achieves a 9.47% higher accuracy than VulDeePecker. SySeVR performs well on the experimental dataset, partly because its set of Common Weakness Enumeration (CWE) vulnerability syntax labels covers most vulnerabilities in the dataset. However, its effectiveness in detecting unknown or real vulnerabilities might be limited. In other words, the coverage of the vulnerability label set is restricted, and vulnerabilities might span multiple lines, making precise localization and slicing challenging. In this experiment, the accuracy is 2.064% lower compared to VulMCI.
Devign utilizes graph neural networks, leveraging node relationships in the graph structure. However, due to its use of complex network node information in the compound graph, it introduces redundant information weakly related to vulnerability nodes, increasing model complexity. Experimental results also show that VulMCI outperforms Devign across various metrics, with an 8.315% increase in True Negative Rate (TNR) and a 7.415% improvement in accuracy. While Devign can identify new vulnerabilities to some extent, it may face challenges when handling large-scale code repositories. VulCNN, using centrality analysis, transforms time-consuming graph analysis into efficient image scanning. It effectively addresses large-scale scanning, but ignoring the discontinuity of pixel information in images results in insufficient vulnerability feature extraction. VulMCI retains richer code relationships in the Code Property Graph (CPG), leading to a 2.877% improvement in True Positive Rate (TPR), a 5.91% increase in TNR, and a 5.162% improvement in accuracy.
While Devign can identify new vulnerabilities to some extent, it may face challenges in handling large-scale code repositories. VulCNN transforms time-consuming graph analysis into efficient image scanning through centrality analysis. However, due to redundant information in the generated image and the discontinuity of pixel information, VulMCI improves upon VulCNN’s weaknesses. VulMCI achieves a 19.601% increase in TPR, a 5.446% increase in TNR, and a 9.231% increase in accuracy, while absorbing its advantages.
3.2.3. Experiments for Answering RQ3
@@
To assess the real-world vulnerability detection capability of VulMCI, we conducted model training and testing on 2011 real vulnerabilities and their corresponding patched versions from the National Vulnerability Database (NVD), maintaining the same data partitioning strategy as in previous experiments. Due to the scarcity of real-world vulnerability data, and to mitigate the potential impact of individual sample variations on result accuracy, we performed five repetitions of the experiments. We saved the results of the final generation of the model, conducted testing on each fold of the test set, calculated the average for each fold, and ultimately obtained the results from five model testing sessions.
To precisely evaluate the model’s generalization ability, we computed the average accuracy and accuracy standard deviation for each experiment. The stability of the experimental results was visually presented using an error bar graph, as depicted in Figure 6. The error bar graph vividly illustrates the average accuracy of each experiment and expresses the dispersion of accuracy in the form of error bars. The length of the error bars provides insights into the fluctuation in model performance across different experiments. Shorter error bars indicate relatively stable model performance, while longer error bars signify higher instability in performance.
The experimental outcomes reveal that VulMCI exhibits strong detection performance on real-world datasets. The average accuracy across five repeated experiments reached 86.88%, with minimal error fluctuations. The model’s performance varied within the range of 0.77% to 0.83%, indicating its adaptability to diverse data distributions. These findings not only offer robust support for practical applications but also suggest promising prospects for real-world deployment of VulMCI.
4. Discussion
The DISCUSSION section highlights key aspects of the entire system process and outlines future directions for development:
4.1. Limitations in Static Analysis Efficiency
@@
Throughout the entire system workflow, the computational cost of static analysis for generating the Code Property Graph (CPG) is a major bottleneck, requiring over twenty hours for a dataset with twenty thousand function samples. We are currently contemplating the reconstruction of a new, lightweight code analysis tool to enhance operational efficiency.
4.2. Pixel Row Oversampling Focus
While enhancing code line continuity, we currently overlook pixel row internal continuity. Ongoing research aims to validate the effectiveness of this concept.
4.3. Future Expansion and Scalability
@@ 5. RELATED WORK
In recent years, researchers have proposed various methods for vulnerability detection, categorizing static detection techniques into code similarity-based and pattern-based methods. Code similarity-based approaches, such as (Kim et al., 2017; Li and Ernst, 2012; Pham et al., 2010; Sajnani et al., 2016), are generally suitable for identifying code clones or repetitions but exhibit poor performance in detecting unknown vulnerabilities.
@@
Pattern-based methods encompass several directions. Rule-based methods rely on human experts to manually construct vulnerability features. Tools like Checkmarx(Checkmarx, 2021), FlawFinder(FlawFinder, 2021), and RATS(Rough Audit Tool for Security, 2021) using this approach often struggle to ensure coverage in practical detection, requiring a high level of expertise and thus yielding suboptimal results. Early machine-learning-based vulnerability detection methods, such as those based on metric program representations, typically employed feature engineering to manually extract vulnerability-related features. Classic methods include code churn(Perl et al., 2015; Zimmermann et al., 2010; Morrison et al., 2015; Meneely et al., 2013; Shin et al., 2011; Shin and Williams, 2013; Walden et al., 2014), code complexity(Zimmermann et al., 2010; Moshtari et al., 2013; Morrison et al., 2015; Younis et al., 2016; Shin et al., 2011; Shin and Williams, 2013; Walden et al., 2014), coverage(Zimmermann et al., 2010; Morrison et al., 2015), dependency(Zimmermann et al., 2010; Morrison et al., 2015), organizational(Zimmermann et al., 2010; Morrison et al., 2015), and developer activity(Perl et al., 2015; Bosu et al., 2014; Meneely et al., 2013; Shin et al., 2011; Meneely and Williams, 2010). In recent years, researchers have started using deep learning methods to automatically extract vulnerability features.
@@
Due to different program representations at various compilation stages, deep learning-based program representation methods can be categorized into three types based on the organizational form of code representation: sequence-based(Grieco et al., 2016; Wu et al., 2017; Russell et al., 2018a; Yan et al., 2021; Li et al., 2021a; Tian et al., 2020; Li et al., 2021c), syntax tree-based(Li et al., 2019; Tanwar et al., 2021; Zhang et al., 2021; Wang et al., 2016; Lin et al., 2017), and graph-based(Duan et al., 2019b, 2020; Cao et al., 2020; Li et al., 2021b; Cheng et al., 2021; Zheng et al., 2021; Wang et al., 2020; Okun et al., 2013). VulDeePecker(Li et al., 2018) slices the program, collects code snippets, and transforms them into corresponding vector representations. Ultimately, these vectors are used to train a Bidirectional Long Short-Term Memory (BLSTM) model for vulnerability detection. SySeVR(Li et al., 2022) first performs syntactic analysis and then semantic slicing on vulnerable code, transforming sliced code into fixed-length vectors fed into a Bidirectional Gated Recurrent Unit (BGRU) model for training. VulDeeLocator(Li et al., 2021c) utilizes intermediate code to define program slices for vulnerability detection, proposing a new variant of Bidirectional Recurrent Neural Network (BRNN), namely BRNN-vdl, for vulnerability detection and localization. Devign(Zhou et al., 2019) is a graph neural network-based model that encodes the original function code into a joint graph structure containing rich program semantics. The gated graph recurrent layer learns node features by aggregating and propagating information in the graph, and the Conv module extracts meaningful node representations for graph-level predictions. VulCNN(Wu et al., 2022) is a scalable graph-based vulnerability detection system that effectively converts function source code into images while preserving program semantics, suitable for large-scale vulnerability scanning.
6. CONCLUSION
In this paper, we propose and validate the effectiveness of high-quality images with strong continuity in improving the detection accuracy of Convolutional Neural Networks (CNNs). Introducing the VulMCI system method as an improvement over the baseline approach of transforming source code into images, we address the issue of color discontinuity. Within VulMCI, an adaptive stitching pixel-row oversampling method is designed to generate more continuous code features. Theoretical analysis and parameter studies are conducted, showcasing high detection performance on the SARD and NVD datasets. This multidimensional approach substantiates the validity of the experimental results.

### v2

@@ IV-A Experiment Settings
IV-B Experimental Results
IV-B1 Experiments for Answering RQ1
To validate the effectiveness of the CFG-based pixel row oversampling method in VulMCI, we additionally designed several oversampling methods for comparison. The specific methods are as follows:
1.
To determine the influence of oversampling iterations on the results, we designed a method based on the K-nearest neighbor (KNN) adjacent line code splicing according to the semantic properties predicted by sent2vec. This method adopts segmented combination splicing to generate new lines of code. Given two adjacent lines of code, Line and Line (where represents the line number), adaptive splicing is performed between the two adjacent lines of code. The splicing process is executed according to the following steps:
From right to left in Line , we take the number obtained from Equation 1 as the length of the substring, and then obtain the substring .
From left to right in Line , we take the number obtained from Equation 2 as the length of the substring, and then obtain the substring .
The new concatenated line is obtained as , and we complete the words at the character level.
Where , k-1 new concatenated lines will be gradually inserted between adjacent lines, where k=1 indicates no splicing, k=2 indicates splicing once, k=3 indicates splicing twice, and so on.
TABLE III: Parameter analysis of k-valued adjacent code concatenation method k Remark FPR FNR Pr Re F1 ACC 1 Baseline without splicing 17.976 8.932 92.225 93.412 91.068 88.687 2 Splicing and inserting 1 line 16.429 8.762 92.577 93.955 91.238 89.22 3 Splicing and inserting 2 lines 18.69 8.252 92.476 93.215 91.748 89 4 Splicing and inserting 3 lines 19.762 8.124 92.367 92.863 91.876 88.812 5 Splicing and inserting 4 lines 17.738 8.72 92.381 93.508 91.28 88.906 6 Splicing and inserting 5 lines 21.19 7.359 92.543 92.445 92.641 89 7 Splicing and inserting 6 lines 20.476 7.954 92.341 92.637 92.046 88.75 VulMCI_CFG_k=2 Splicing only CFG nodes, using k=2 splicing strategy 1.931 3.383 97.91 99.237 96.617 97.02 The experimental results, as depicted in TableIII, encompassed a total of 7 scenarios. From these experiments, it was observed that the best performance was achieved when , indicating that inserting a single line yielded the optimal outcome. This approach resulted in a 0.533% increase in accuracy compared to not splicing lines. However, the accuracy was still 7.8% lower compared to VulMCI’s CFG-based method. This discrepancy can be attributed to the indiscriminate oversampling of all lines using the K-nearest neighbor splicing method. While this approach enhanced the continuity of the images, it also fused redundant code segments, including those weakly associated with vulnerabilities. As vulnerability-related code segments are in the minority, their significance was not significantly bolstered by the splicing process.
2.
To assess the impact of different oversampling methods on the results, we devised three additional control methods, including the complete concatenation of adjacent code and an oversampling approach where code lines are first vectorized and then inserted by adding their respective vectors. The complete concatenation of adjacent code involves inserting a single line between adjacent lines, where concatenated_code = left_code + right_code. The oversampling method of inserting vectors involves the insertion vector being equal to the sum of the left and right vectors. To prevent excessively large values resulting from vector addition, we also experimented with dividing the sum of the vector rows by two before insertion, denoted as Insertion_vector = (left_vector + right_vector) / 2.
TABLE IV: Comparison of other splicing methods Method Description FPR FNR F1 Pr Re ACC No splicing Baseline 17.976 8.932 92.225 93.412 91.068 88.687 All Complete code concatenation of adjacent lines 17.5 8.464 92.559 93.606 91.536 89.157 Vec Vectorize first, then add adjacent vector rows 19.524 8.379 92.268 92.925 91.621 88.687 Vec2 Divide the sum of adjacent vector rows by two before insertion 22.857 7.529 92.177 91.885 92.471 88.436 VulMCI_CFG_All Only splice CFG nodes and adopt the splicing strategy ”All” 1.609 3.259 98.035 99.364 96.741 97.199 The experimental results, as shown in TableIV, indicate that the methods directly processing vector rows (vec and vec2) perform poorly, with vec2 even performing slightly worse than no splicing, suggesting that directly oversampling vector rows may lead to information loss or confusion. The methods of complete code concatenation and adaptive concatenation of 1 line (k=2) yield similar results, hence we compared them with the CFG-based method of adjacent lines with complete concatenation (VulMCI_CFG_all), which shows slightly better performance than VulMCI_CFG_k=2. This difference in results may be related to the distribution of the concatenated data.
3.
As the Code Property Graph (CPG) contains various types of relationship subgraphs, to investigate which type of edge concatenation performs best in experiments, we compared three types of concatenation: Data Dependency Graph (DDG), Control Dependency Graph (CDG), and Control Flow Graph (CFG). Since in the above experiments, the concatenation methods ’k=2’ and ’all’ performed the best and yielded similar results, we also applied two different concatenation methods to the three types of edges.
TABLE V: Comparison of other splicing methods Method Description FPR FNR F1 Pr Re ACC No Concatenation Baseline 17.976 8.932 92.225 93.412 91.068 88.687 Concatenate All Concatenate DDG, CFG, and CDG, complete code 17.619 8.55 92.493 93.56 91.45 89.063 DDG_k=2 Concatenate only DDG nodes, split-combine approach 3.004 3.053 97.876 98.823 96.947 96.961 CDG_k=2 Concatenate only CDG nodes, split-combine approach 15.88 11.056 91.201 93.576 88.944 87.604 CFG_k=2 Concatenate only CFG nodes, split-combine approach 1.931 3.383 97.91 99.237 96.617 97.02 DDG_all Concatenate only DDG nodes, Complete splice approach 2.575 3.012 97.979 98.989 96.988 97.11 CDG_all Concatenate only CDG nodes, Complete splice approach 12.768 10.52 92.063 94.799 89.48 88.856 CFG_all Concatenate only CFG nodes, Complete splice approach 1.609 3.259 98.035 99.364 96.741 97.199 The experimental results, as shown in Table 5, indicate that when concatenating based on edge relationships, concatenating complete adjacent lines yields better results than using split-combine concatenation. This is related to the data distribution. When all lines of code are concatenated indiscriminately, the distances between lines are closer, making split-combine concatenation more suitable for this data distribution. The sent2vec model can predict the concatenated lines well by connecting contextual semantics. However, when concatenating based on code structural relationships, the concatenated code is more directional and strongly correlated with the vulnerability execution process, resulting in cases of concatenation across lines. Therefore, concatenating complete adjacent lines based on code structural information is more suitable. Among the three edge types, concatenating based on CFG yields the best results, with an F1 score of 98.035% and an accuracy of 97.199%. This is because CFG can better capture code structure and control flow. Specifically, code execution follows temporal logic, and concatenating based on the dependencies of data flow and code control flow does not disrupt the temporal relationship of the code. It also narrows the distance between key statements, facilitating vulnerability feature extraction. Concatenating based on DDG yields the second-best results because DDG only focuses on vulnerabilities caused by data, such as buffer overflow, and does not well reflect the formation relationships of logical vulnerabilities. Concatenating based on CDG yields the worst results, although slightly better than no concatenation. This is because CDG generally only contains conditional branches, such as if and while. Finally, when concatenating all three edge types of DDG, CFG, and CDG, the accuracy is only 0.376% higher than no concatenation. It is evident that simultaneously concatenating three types of edges leads to many repeated concatenations between nodes, resulting in a loss of distinction between key vulnerability statements, thus having a counterproductive effect.
IV-B2 Experiments for Answering RQ2
@@
In this section, we compare VulMCI with several vulnerability detection tools, including a commercial static vulnerability detection tool (i.e., Checkmarx[9]), two open-source static analysis tools (i.e., FlawFinde[7] and RATS[8]), and five deep learning-based vulnerability detection methods (i.e., VulDeePecker[10], SySeVR[15], Devign[13], VulCNN[14] and AMPLE[23]). Figure 6 presents a bar chart that intuitively illustrates the metrics of True Positive Rate (TPR), True Negative Rate (TNR), and Accuracy (ACC) for each of these methods.
@@
For the three static analysis tools, RATS, Checkmarx, and FlawFinder, their detection effectiveness is not satisfactory, with accuracy rates all below 60%. One plausible explanation is that they rely on human experts to define vulnerability rules for detection. However, as vulnerability types become increasingly complex, human experts cannot comprehensively define patterns for all vulnerabilities, leading to higher rates of false positives and false negatives.
@@
Regarding VulDeePecker and SySeVR, both methods employ program slicing to process datasets, vectorizing slices for training bidirectional recurrent neural networks (BRNN) for vulnerability detection. The performance disparity between these two systems arises from VulDeePecker only considering vulnerabilities caused by API function calls and solely focusing on data dependency information. In contrast, SySeVR considers semantic information resulting from both data and control dependencies and supports a wider range of vulnerability syntax labels for matching vulnerabilities. Consequently, SySeVR achieves a 9.47% higher accuracy than VulDeePecker.
SySeVR utilizes code slicing for vulnerability feature extraction, leading to incomplete code semantics. On the other hand, VulMCI enhances the semantic information of critical code while retaining complete function code. Additionally, the limited coverage of vulnerability label sets in slicing methods means that a function sample may have multiple vulnerability candidates, making precise vulnerability localization and slicing challenging. VulMCI achieves a 10.70% higher accuracy than SySeVR.
Devign employs a graph neural network approach, effectively leveraging node relationships within the graph structure. However, its use of compound graphs with intricate network node information introduces redundant information weakly correlated with vulnerability nodes, increasing model complexity and performance overhead. Moreover, as the number of nodes increases, the model struggles to learn edge node information, leading to information loss. In contrast, VulMCI utilizes an image-based approach for detection, enhancing vulnerability features through oversampling based on code structure while retaining complete code semantics, thus avoiding the aforementioned issues.
AMPLE addresses Devign’s shortcomings by simplifying nodes according to type and variable, representing different types of edges with weighted vectors, and enhancing node representation using a multi-head attention mechanism. This results in a 6.41% increase in accuracy compared to Devign. However, due to the addition of modules, AMPLE still incurs considerable performance overhead. In contrast, VulMCI achieves a higher simplification rate by mapping the complex node network to a network between line numbers in the preprocessing phase. Experimental results demonstrate that VulMCI outperforms AMPLE by 3.60% in accuracy.
VulCNN utilizes centrality analysis to transform time-consuming graph analysis into efficient image scanning, effective for large-scale scanning. However, it overlooks the discontinuity of pixel information in images, leading to insufficient vulnerability feature extraction. VulMCI addresses this issue through pixel row oversampling, resulting in a 19.82% increase in true positive rate (TPR), a 4.95% increase in true negative rate (TNR), and an 8.89% increase in accuracy (ACC).
IV-B3 Experiments for Answering RQ3
@@
To evaluate VulMCI’s real-world vulnerability detection capabilities, we trained and tested the model using 2011 real vulnerabilities and their corresponding fixed version files from the NVD. The data partitioning method remained consistent with previous experiments. Due to the scarcity of real-world vulnerability data and to prevent inaccuracies caused by individual sample variations, we conducted five repeated experiments, saving the results of the final generation. Each fold of the test set was evaluated in each experiment, and the average result of each fold’s testing was computed. The average F1 score over the five repeated experiments reached 89.07%, with an average accuracy of 90.41%.
Real-world vulnerability detection often involves large-scale code scanning, necessitating consideration of performance overhead. The time overhead of each system is depicted in Figure 8, with VulCNN exhibiting excellent performance in this regard. We augmented VulCNN with a pixel row oversampling module to reduce the computation of the three central indicators for the PDG and trained it using grayscale images. Meanwhile, VulMCI also selected 100 lines as the function length threshold without increasing model training overhead. Figure 7 illustrates the time overhead proportions of the six models, among which VulMCI’s time overhead is similar to VulCNN’s, approximately one-sixth of Devign’s, and approximately one-fourth of SySeVR’s.
V Discussion
@@
Throughout the entire system workflow, the static analysis for generating the Code Property Graph (CPG) incurs the highest computational cost. Processing a dataset comprising twenty thousand function samples requires over twenty hours. Additionally, there is some redundancy in the label information of analyzed nodes. We are currently contemplating the reconstruction of a new, lightweight code analysis tool to enhance operational efficiency.
While utilizing the pixel row oversampling method to generate more continuous code line features, we have focused solely on oversampling between code lines. In other words, we have considered only the continuity of pixel rows above and below the code lines, neglecting the continuity within pixel rows. However, we are committed to ongoing research to substantiate the effectiveness of this concept.
@@ VI RELATED WORK
In recent years, researchers have proposed various methods for vulnerability detection, categorizing static detection techniques into code similarity-based and pattern-based methods. Code similarity-based approaches, such as [3, 4, 6, 24], are generally suitable for identifying code clones or repetitions but exhibit poor performance in detecting unknown vulnerabilities.
@@
Pattern-based methods encompass several directions. Rule-based methods rely on human experts to manually construct vulnerability features. Tools like Checkmarx[9], FlawFinder[7], and RATS[8] using this approach often struggle to ensure coverage in practical detection, requiring a high level of expertise and thus yielding suboptimal results. Early machine-learning-based vulnerability detection methods, such as those based on metric program representations, typically employed feature engineering to manually extract vulnerability-related features. Classic methods include code churn[25, 26, 27, 28, 29, 30, 31], code complexity[26, 32, 27, 33, 29, 30, 31], coverage[26, 27], dependency[26, 27], organizational[26, 27], and developer activity[25, 34, 28, 29, 35]. In recent years, researchers have started using deep learning methods to automatically extract vulnerability features.
@@
Due to different program representations at various compilation stages, deep learning-based program representation methods can be categorized into three types based on the organizational form of code representation: sequence-based[36, 37, 38, 39, 40, 41, 42], syntax tree-based[43, 44, 45, 46, 47], and graph-based[48, 49, 50, 51, 52, 53, 54, 55]. VulDeePecker[10] slices the program, collects code snippets, and transforms them into corresponding vector representations. Ultimately, these vectors are used to train a Bidirectional Long Short-Term Memory (BLSTM) model for vulnerability detection. SySeVR[15] first performs syntactic analysis and then semantic slicing on vulnerable code, transforming sliced code into fixed-length vectors fed into a Bidirectional Gated Recurrent Unit (BGRU) model for training. VulDeeLocator[42] utilizes intermediate code to define program slices for vulnerability detection, proposing a new variant of Bidirectional Recurrent Neural Network (BRNN), namely BRNN-vdl, for vulnerability detection and localization. Devign[13] is a graph neural network-based model that encodes the original function code into a joint graph structure containing rich program semantics. The gated graph recurrent layer learns node features by aggregating and propagating information in the graph, and the Conv module extracts meaningful node representations for graph-level predictions. AMPLE[23] simplified node representations based on Devign and designed an edge-aware graph convolutional network module, addressing the challenge of graph neural networks capturing relationships between distant graph nodes. VulCNN[14] is a scalable graph-based vulnerability detection system that effectively converts function source code into images while preserving program semantics, suitable for large-scale vulnerability scanning.
VII CONCLUSION
In this paper, we propose and validate that high-quality images with strong continuity can effectively improve the detection accuracy of Convolutional Neural Networks (CNNs). Additionally, we introduce the use of CFG-based pixel row oversampling method, which synthesizes new samples by concatenating between CFG node code, enhancing the continuity of code images. We conduct theoretical analysis and parameter studies, design and compare multiple oversampling methods, as well as some advanced vulnerability detection systems. VulMCI achieves high detection performance and low time overhead in both the SARD and NVD datasets, demonstrating its effectiveness.